# Four loop renormalization of $3$-quark operators in QCD

### v1

@@ 2 General operator
Having established the mapping of the evanescent $\mathbb{C}_{qrs}$ structures to the four dimensional ones these are applied to the evaluation of the Greens’s function before the renormalization constant for ${\cal O}^{ijk}_{\alpha\beta\gamma}$ is extracted in the $\overline{\mbox{MS}}$ scheme [12, 13]. In other words the $\epsilon$ expansion of the $d$-dependent coefficients of each term in (3.6) is effected first. This results in the four dimensional anomalous dimension
@@
numerically. It is worth recalling that the $\mathbb{C}_{qrs}$ structures that remain are, [12, 13],
@@
$\left.\frac{}{}\mathbb{C}_{220}\right|_{d=4}$ $=$ $\left[\sigma^{\mu\nu}\otimes\sigma_{\mu\nu}\otimes{\cal I}~+~\sigma^{\mu\nu}\otimes{\cal I}\otimes\sigma_{\mu\nu}~+~{\cal I}\otimes\sigma^{\mu\nu}\otimes\sigma_{\mu\nu}\right]$
@@
where $\Delta^{2}$ == 0. The right and left handed quarks are denoted by $\psi_{R}$ == $\mbox{\small{$\frac{1}{2}$}}(1+\gamma^{5})\psi$ and $\psi_{L}$ == $\mbox{\small{$\frac{1}{2}$}}(1-\gamma^{5})\psi$ respectively, the first superscript represents the spin that is not part of the bilinear part of the operator and the second relates to the bilinear properties itself, and the subscript sign indicates the chirality [12, 20]. For more details on the discrete and continuous symmetries of these operators see the discussion in [20]. To extract the respective anomalous dimensions merely requires substituting the eigenvalues of the respective $\mathbb{C}_{qrs}$ structures that were deduced in [12, 20] and recorded in Table 1. Consequently we have
@@
where we reproduced the lower loop results of [11, 12, 13]. This provides a minor check since [13] did not use Forcer. Numerically we have
@@
Acknowledgement. The author thanks the DESY Zeuthen theory group, where part of this work was carried out, for their hospitality and Kolleg Mathematik Physik Berlin for a Visiting Scholarship. The research was carried out with the support of the STFC Consolidated Grant ST/X000699/1 and was undertaken on Barkla, part of the High Performance Computing facilities at the University of Liverpool, UK. For the purpose of open access, the author has applied a Creative Commons Attribution (CC-BY) licence to any Author Accepted Manuscript version arising.

### v2

@@ 2 General operator
Having established the mapping of the evanescent $\mathbb{C}_{qrs}$ structures to the four dimensional ones these are applied to the evaluation of the Green’s function before the renormalization constant for ${\cal O}^{ijk}_{\alpha\beta\gamma}$ is extracted in the $\overline{\mbox{MS}}$ scheme [12, 13]. In other words the $\epsilon$ expansion of the $d$-dependent coefficients of each term in (3.6) is effected first. This results in the four dimensional anomalous dimension
@@
numerically. One observation deserves comment and it concerns the comparison of the results of (3.2) with (3.9). If one formally replaces the $\mathbb{C}_{qrs}$ matrices in (3.2) with their strictly four dimensional counterparts then (3.9) is reproduced. By this we mean the evanescent matrices of (3.2) are replaced using the relations of (3.5) and (3.6) but with $d$ replaced by $4$. This is consistent with the general formal relation
$\gamma_{\cal O}(a)~=~\lim_{d\to 4}\gamma_{\cal O}^{\mbox{\scriptsize{naive}}}(a)$ (3.11)
which additionally provides a check on the translation of the two actual renormalization constants into their respective $\overline{\mbox{MS}}$ anomalous dimensions. It is worth recalling that the $\mathbb{C}_{qrs}$ structures that remain in (3.9) are, [12, 13],
@@
$\left.\frac{}{}\mathbb{C}_{220}\right|_{d=4}$ $=$ $\sigma^{\mu\nu}\otimes\sigma_{\mu\nu}\otimes{\cal I}~+~\sigma^{\mu\nu}\otimes{\cal I}\otimes\sigma_{\mu\nu}~+~{\cal I}\otimes\sigma^{\mu\nu}\otimes\sigma_{\mu\nu}$
@@
where $\Delta^{2}$ == 0. The right and left handed quarks are denoted by $\psi_{R}$ == $\mbox{\small{$\frac{1}{2}$}}(1+\gamma^{5})\psi$ and $\psi_{L}$ == $\mbox{\small{$\frac{1}{2}$}}(1-\gamma^{5})\psi$ respectively, the first superscript represents the spin that is not part of the bilinear part of the operator and the second relates to the bilinear properties itself, and the subscript sign indicates the chirality [12, 20]. For more details on the discrete and continuous symmetries of these operators see the discussion in [20]. To extract the respective anomalous dimensions merely requires substituting the eigenvalues of the respective $\mathbb{C}_{qrs}$ structures that were deduced in [12, 20, 35] and recorded in Table 1. Consequently we have
@@
where we reproduced the lower loop results of [11, 12, 13] which represents a minor check since [13] did not use Forcer. Another check is provided by the large $N_{\!f}$ result of [36] where the coefficients of the leading order $N_{\!f}$ terms of $\gamma^{(\mbox{\small{$\frac{1}{2}$}},0)}_{\pm}(a)$ were computed. It is therefore reassuring to note the coefficient of $N_{\!f}^{3}$ in the four loop term of both these anomalous dimensions is in agreement with the second expression of equation (21) in [36]. Numerically we have
@@
Acknowledgement. The author thanks the DESY Zeuthen theory group, where part of this work was carried out, for their hospitality and Kolleg Mathematik Physik Berlin for a Visiting Scholarship. We thank L. Vecchi for correspondence. The research was carried out with the support of the STFC Consolidated Grant ST/X000699/1 and was undertaken on Barkla, part of the High Performance Computing facilities at the University of Liverpool, UK. For the purpose of open access, the author has applied a Creative Commons Attribution (CC-BY) licence to any Author Accepted Manuscript version arising.